A solution is prepared by dissolving 38.7g of nonelectrolyte into 218g of water.


The freezing point of the solution is measured to be -5.53°C. Calculate the


molar mass of the solute. kf of water = 1.86°C/m

1
Expert's answer
2021-11-18T09:44:03-0500

Freezing point depression= Freezing point of water-Freezing point of solution

= 0-(-5.53)= 5.53 °C

Freezing point depression= Freezing point constant of water "\times" molality

molality = Freezing point depression/Freezing point constant of water

molality "= \frac{5.53}{1.86}= 2.973 \; m"

molaity=moles of solute/mass of water in kg

moles of solute= molality "\times" mass of water in kg

moles of solute "= 2.973 \times 0.218= 0.648" moles of solute

molar mass of solute "= \frac{mass}{moles} = \frac{38.7}{0.648}= 59.7 \;g/mol = 60 \; g/mol"